OpenStudy (anonymous):

75% of what number is 66?

OpenStudy (anonymous):

88

OpenStudy (king):

ya 88

OpenStudy (anonymous):

75% = 3/4 3/4= 66 thus, 1/4= 22 4/4=88

OpenStudy (anonymous):

88